George needs to earn a B in his Algebra class. His current test scores are 88,81,78, and 85. His final exam is worth 4 test scores. In order to earn a B George's average must lie between 80 and 89 inclusive. What range of scores can George receive on the final exam to earn a B in the course
Thank you

Answer by swincher4391(1107)

Assume that his final grade is worth 8 test scores.
Then he needs to have 80*8 = 640 through 89*8 = 712 total points.
614<= 88+81+78+85+4x<=712
282<=4x<= 380
70.5<=x<=95